1. If the weight is descending freely, determine the tension in the cord both before and after a brake force P – 100 lb is applied. Neglect the thickness of brake.


1
Expert's answer
2021-07-30T16:37:01-0400

Taking mass moment inertia

"\sum M_o=I_o\alpha \implies200*2-F_B(3)=80lb"

"and F_B=f_s*N_B=f_s*(4p)"

"A_s=p=100lb\implies F_B=80lb" and "\alpha =13.33rad ^{-1}"

"\sum f_y=0 \implies F_B+200=T \implies 280lb"

"T_i=F_Be^{t_i*Q}= T_i=80*e^{0.20*Q_1}"

"T_2=80*e^{0.20*Q_2}\\
Q_2=Tan(71.6^0)\\
Q_1=tan(-45^0)\\
\therefore T_1=146lb ; T_2=65.2lb"